# SynthFormer: Equivariant Pharmacophore-based Generation of Molecules for Ligand-Based Drug Design

### v1

@@ 5 Conclusions
The evaluation of the generative model demonstrates its capacity to generate chemically diverse and relevant compounds while maintaining desirable physicochemical properties. By employing a variety of similarity metrics—such as Tanimoto, Murcko, and Gobbi fingerprints—and analyzing structural properties, the generated molecules are shown to maintain low similarity to reference ligands, while preserving pharmacophore information. Furthermore, the genetic algorithm-based optimization effectively guides the exploration of new chemical spaces, improving molecular properties while preserving scaffold integrity. The embedding analysis further validates the model’s ability to capture key molecular features, such as ring structures and nitrogen atoms, illustrating its proficiency in maintaining meaningful chemical relationships. These results indicate that SynthFormer has strong potential to generate novel, drug-like molecules that can support hit identification, hit expansion and lead optimization efforts in drug discovery.
Appendix A Appendix
Appendix B Generated active ligands
Figure 4: The first molecule corresponds to the reference structure derived from the 3COY PDB ID. The subsequent molecules are computationally generated using the Synthformer model, illustrating its capability to design novel molecular structures based on a known protein-ligand complex. These results highlight Synthformer’s potential in generating diverse and plausible molecular candidates.
Figure 5: The first molecule corresponds to the reference structure derived from the 3COY PDB ID. The subsequent molecules are expanded hits generated by the Synthformer model, demonstrating structural diversity and novel chemical scaffolds.

### v2

@@ 5 Conclusions
SynthFormer presents a novel approach in drug discovery by effectively generating synthetically accessible molecules directly from pharmacophores. Through its integration of a 3D equivariant GNN and a synthesis-aware decoding mechanism, it addresses the critical challenge of designing both active and synthesizable molecules. The results demonstrate that SynthFormer not only outperforms random baselines in docking performance but also achieves higher synthesisability compared to existing 3D models. These findings highlight the model’s potential to accelerate drug discovery by optimizing molecular design and exploring the synthesizable chemical space around hit compounds, marking it as a promising tool for advancing the field. Moreover, it has shown effectiveness in hit expansion and molecule optimization, further expanding its applicability in drug discovery workflows.